# Physics-Informed LLM-Agent for Automated Modulation Design in Power Electronics Systems

### v1

@@ 5.2. Experimental Results
Table 3. Experimental results with training / validation / test set split of 5% (10 samples), 10% (20 samples) and 85% (170 samples), respectively. LP-COMDA statistically outperforms baselines (marked with *), with a 63.2% lower error as compared to the second-best benchmark (TST),measured in Mean Absolute Errors (MAE).
Model Training Validation Test p-value
BN 2.455$\pm$0.0 2.206$\pm$0.0 2.348$\pm$0.0 5.03E-14*
SVR 2.049$\pm$0.0 1.780$\pm$0.0 1.991$\pm$0.0 1.25E-13*
XGBoost 1.321$\pm$0.0 2.194$\pm$0.0 1.986$\pm$0.0 1.27E-13*
RF 3.167$\pm$0.0 3.937$\pm$0.0 3.622$\pm$0.0 4.89E-15*
LSTM 1.246$\pm$0.165 1.129$\pm$0.078 1.193$\pm$0.098 1.89E-6*
GRU 1.196$\pm$0.127 1.356$\pm$0.134 1.270$\pm$0.114 2.80E-6*
LN-GRU 0.534$\pm$0.097 0.480$\pm$0.075 0.527$\pm$0.086 2.33E-4*
TCN 0.685$\pm$0.018 0.744$\pm$0.062 0.726$\pm$0.061 3.79E-6*
TST 0.264$\pm$0.033 0.240$\pm$0.033 0.264$\pm$0.031 3.93E-3*
TSiTPlus 0.585$\pm$0.099 0.567$\pm$0.082 0.569$\pm$0.099 2.61E-4*
MiniRocket 0.522$\pm$0.305 0.616$\pm$0.346 0.646$\pm$0.343 0.0246*
GRU-VAE 0.686$\pm$0.074 0.663$\pm$0.036 0.721$\pm$0.056 2.42E-6*
CirNet Only 0.162$\pm$0.032 0.239$\pm$0.014 0.234$\pm$0.007 3.98E-3*
LP-COMDA 0.170$\pm$0.013 0.191$\pm$0.005 0.201$\pm$0.006 N/A
@@
Table 4. Experimental results with training / validation / test set split of 50% (100 samples), 10% (20 samples) and 40% (80 samples), respectively. LP-COMDA statistically outperforms baselines (marked with *), with a 23.7% lower error as compared to the second-best benchmark (TST), measured in Mean Absolute Errors (MAE).
@@
Table 3 and 4 summarise the experimental results for LP-COMDA and baseline for two different splits. Each algorithm with the respective optimal hyperparameters was ran for ten times. The average MAE for training, validation and test sets are reported. The reported P-values are the results of T-tests at a significance level of 0.02.
@@
For both splits, experimental results show that LP-COMDA outperforms conventional machine learning approaches, such as BN, SVR, XGBoost and RF, and exhibits superior performance over deep learning predictive models, such as LSTM, GRU, LN-GRU and TCN, and predictive models such as GRU-VAE. When compared to state-of-the-art models like TST, TSiTPlus, and MiniRocket, LP-COMDA again demonstrates notably enhanced forecasting accuracy. To further affirm the hierarchical PINNs structure, we contrasted CirNet with LP-COMDA. Noticeably, LP-COMDA with 10 training samples outperforms most benchmarks with 100 training samples, and align with CirNet-Only model.
@@
To visualize the modeling performance, Figure 5 presents the modeled $i_{L}$ waveform for both LP-COMDA and benchmark models. Notably, the LP-COMDA modeling outcomes align more congruently with the measurements than any of the other models, demonstrating its outstanding modeling precision.
@@
Overall, we conclude that the proposed LP-COMDA statistically achieves better modeling results than other baseline models for power electronics modeling, even with extremely low-data scenario.
@@
1
To define a modulation
strategy $S$ to design
I need your help to design my
dual active bridge converter.
Sure, I can help you with that. let’s start with choosing the
modulation strategy. Which one would you like to use for
your dual active bridge converter, TPS, DPS or Hybrid?
2
To define the prioritized
objective $O$
TPS.
Great! Now let’s move on to the prioritized optimization
objective. Which of the following objectives would you like
to prioritize: Efficiency, Soft switching, or Current stress?
3
To define operation
conditions
Current stress.
Understood.Let’s discuss the operating conditions of your
dual active bridge converter. Could you provide the following
information: rated power, rated input voltage, rated output
voltage, actual operating power, actual output voltage.
4
To select optimization
algorithms
Rated power: 1000 W
Rated input voltage: 200 V
Rated output voltage: 200 V
Operating power: 600W
Actual output voltage: 160V
Thank you for providing the operating conditions. Lastly, we
need to choose an optimization algorithm. Which optimization
algorithm would you choose: PSO, DE or GA?
@@
PSO.
OK, I have collected all the parameters and am about to
execute the program.
@@
Table 5. An use case exemplar: with 5 request series, LP-COMDA collects all necessary design specifications, which include a modulation strategy $S$ to design, prioritized objective $O$, operating conditions, and the selected optimization algorithms. After specification collection via text input, the design outcome is presented with text and figures.
6. User Study
To assess LP-COMDA’s empirical performance in improving engineering efficiency, we design a practical use case for the modulation strategy of the DAB converter and conduct an empirical experiment with 20 industrial practitioners.
The user requirements for this design process include:
  • Guidance in defining design specifications.
  • Optimal design outcomes for the tailored scenario.
  • Analysis on the design outcomes.
Based on the use case, we conduct a user study with 20 industrial practitioners recruited via research labs and collaborating external companies on a voluntary basis. The participants consist of 10 junior engineers with less than three years of experience, and 10 senior engineers with over five years of relevant experience. All participants declared to have no prior experience in using LLM to assist engineering design works.
All participants are given two independent variants of the same task, namely to design the TPS modulation strategy for a Dual Active Bridge (DAB) converter to serve as a DC transformer of the DC-DC power grid, as shown in Fig. 1. We ask the participant to first use LP-COMDA’s chat interface to address task variant one. After completion, they then use standard practice (i.e. analytical and manual design) based on the given design requirements and data to address task variant two. To assess the effectiveness, the effective working hours are measured in 30-minute time blocks while tackling the tasks.
Table 5 shows an example of a multi-turn conversation between the user and LP-COMDA, as well as the output components. Empirically, the results of the user study are highly promising, in terms of both functionality and efficiency. In terms of functionality, all 20 participants are able to complete both assigned tasks successfully. In terms of efficiency, the 10 junior engineers spent an average of 1.2 time blocks and 115.5 time blocks in completing the tasks using LP-COMDA and analytical approach, respectively; where the 10 senior engineers recorded an average of 0.9 time blocks and 30.5 time blocks with LP-COMDA and analytical approach, respectively. Overall, we observe that Junior engineers and senior engineers experienced 96.3 times and 33.9 times efficiency, respectively.
7. Conclusion
Modulation strategy design of power converters is pivotal for the optimal functioning of renewable energy power systems. However, numerous AI-centric modulation design automation systems today grapple with challenges tied to explainability, usability, scalability, and data intensity. In our research, we introduced LP-COMDA, a LLM-based Physics-informed autonomous agent to perform effective power converter modulation design automation. Experiments demonstrate LP-COMDA’s proficiency in accurate power converter modeling, even with a data-constrained environment. Our user study shows that using LP-COMDA helps engineers improve efficiency by over 33 times in practice.

### v2

@@ 3 $D_{1}$ $D_{o}$ $D_{2}$
Table 3: Experimental results with training / validation / test set split of 5% (10 samples), 10% (20 samples) and 85% (170 samples), respectively. PHIA statistically outperforms baselines (marked with *), with a 63.2% lower error as compared to the second-best benchmark (TST),measured in Mean Absolute Errors (MAE).
Model
Training Validation Test p-value
[…]
@@
Table 4: Experimental results with training / validation / test set split of 50% (100 samples), 10% (20 samples) and 40% (80 samples), respectively. PHIA statistically outperforms baselines (marked with *), with a 23.7% lower error as compared to the second-best benchmark (TST), measured in Mean Absolute Errors (MAE).
@@
Table 3 and 4 summarise the experimental results for PHIA and baseline for two different splits. Each algorithm with the respective optimal hyperparameters was ran for ten times. The average MAE for training, validation and test sets are reported. The reported P-values are the results of T-tests at a significance level of 0.02.
@@
For both splits, experimental results show that PHIA outperforms conventional machine learning approaches, such as BN, SVR, XGBoost and RF, and exhibits superior performance over deep learning predictive models, such as LSTM, GRU, LN-GRU and TCN, and predictive models such as GRU-VAE. When compared to state-of-the-art models like TST, TSiTPlus, and MiniRocket, PHIA again demonstrates notably enhanced forecasting accuracy. To further affirm the hierarchical PINNs structure, we contrasted CirNet with PHIA. Noticeably, PHIA with 10 training samples outperforms most benchmarks with 100 training samples, and align with CirNet-Only model.
@@
To visualize the modeling performance, Figure 5 presents the modeled $i_{L}$ waveform for both PHIA and benchmark models. Notably, the PHIA modeling outcomes align more congruently with the measurements than any of the other models, demonstrating its outstanding modeling precision.
@@
Overall, we conclude that the proposed PHIA statistically achieves better modeling results than other baseline models for power electronics modeling, even with extremely low-data scenario.
Interpretation
Accuracy and Reliability
The improved accuracy observed in the experiments, particularly in low-data scenarios, underscores the robustness of the proposed framework. This enhancement ensures reliable performance, which is crucial for reducing the risk of failures in real-world applications. The system’s ability to achieve high precision despite limited data sets demonstrates its adaptability and suitability for resource-constrained environments.
Efficiency Gains
The significant reduction in error rates translates directly to efficiency gains. By optimizing workflows, the framework enables engineers to achieve desired outcomes with fewer iterations and reduced computational resources. This efficiency not only accelerates the design process but also minimizes the operational costs associated with traditional methods.
Scalability
Performance improvements across metrics indicate better scalability of the framework. This scalability allows the system to handle more complex systems or broader application domains without requiring proportional increases in computational or human resources. Such scalability is essential for extending the applicability of the framework to diverse use cases.
Explainability and Transparency
The integration of hierarchical modeling enhances the explainability of the results. By aligning the model predictions with physical principles, the framework produces outputs that are both interpretable and trustworthy. This transparency is crucial for fostering confidence in the system’s decisions, especially in critical domains like power electronics.
Implications for Decision-Making
The precision and reliability of the framework support informed decision-making in key applications. For example, in renewable energy systems and power grid operations, the metrics directly influence operational success and system stability. The enhanced performance ensures that the framework can be effectively deployed in high-stakes scenarios.
This interpretation highlights the broad implications of the observed performance improvements, showcasing the framework’s potential to revolutionize practical applications and advance theoretical understanding in the domain.
User Study
To assess PHIA’s empirical performance in improving engineering efficiency, we design a practical use case for the modulation strategy of the DAB converter and conduct an empirical experiment with 20 industrial practitioners.
Specifically, the user requirements for this design process include: (1) Guidance in defining design specifications; (2) Optimal design outcomes for the tailored scenario; and (3) Analysis on the design outcomes.
Based on the use case, we conduct a user study with 20 industrial practitioners recruited via research labs and collaborating external companies on a voluntary basis. The participants consist of 10 junior engineers with less than three years of experience, and 10 senior engineers with over five years of relevant experience. All participants declared to have no prior experience in using AI systems to assist engineering design works.
All participants are given two independent variants of the same task, namely to design the TPS modulation strategy for a Dual Active Bridge (DAB) converter to serve as a DC transformer of the DC-DC power grid, as shown in Fig. 1. We ask the participant to first use PHIA’s chat interface to address task variant one. After completion, they then use standard practice (i.e. analytical and manual design) based on the given design requirements and data to address task variant two. To assess the effectiveness, the effective working hours are measured in 30-minute time blocks while tackling the tasks.
Table 5 (in the Appendix) shows an example of a multi-turn conversation between the user and PHIA, as well as the output components. Empirically, the results of the user study are highly promising, in terms of both functionality and efficiency. In terms of functionality, all 20 participants are able to complete both assigned tasks successfully. In terms of efficiency, the 10 junior engineers spent an average of 1.2 time blocks and 115.5 time blocks in completing the tasks using PHIA and analytical approach, respectively; where the 10 senior engineers recorded an average of 0.9 time blocks and 30.5 time blocks with PHIA and analytical approach, respectively. Overall, we observe that Junior engineers and senior engineers experienced 96.3 times and 33.9 times efficiency, respectively.
Conclusion
Modulation strategy design of power converters is pivotal for the optimal functioning of renewable energy power systems. However, numerous AI-centric modulation design automation systems today grapple with challenges tied to explainability, usability, scalability, and data intensity. In our research, we introduced PHIA, a LLM-based Physics-informed autonomous agent to perform effective power converter modulation design automation. Our physics-informed approach combines the reasoning capabilities of large language models with rigorous physical constraints, enabling robust and interpretable design solutions. Experiments demonstrate PHIA’s proficiency in accurate power converter modeling, even with a data-constrained environment. The hierarchical physics-informed neural network architecture ensures both accuracy and interpretability, while the LLM-based planning module provides an intuitive interface for engineers. Our comprehensive user study shows that using PHIA helps engineers improve efficiency by over 33 times in practice, highlighting its potential for widespread adoption in industrial applications. Future work could explore extending PHIA to other power electronics applications and incorporating additional physics-based constraints to further enhance its capabilities.
Acknowledgement
This research is supported in part by the Ministry of Education, Singapore, under its Academic Research Fund Tier 2 (Award No. MOE-T2EP20123-0015). Any opinions, findings and conclusions, or recommendations expressed in this material are those of the authors and do not reflect the views of the Ministry of Education, Singapore.
Appendix A Appendix A: Example Use Case
@@
1 To define a modulation strategy $S$ to design I need your help to design my dual active bridge converter. Sure, I can help you with that. let’s start with choosing the modulation strategy. Which one would you like to use for your dual active bridge converter, TPS, DPS or Hybrid?
2 To define the prioritized objective $O$ TPS. Great! Now let’s move on to the prioritized optimization objective. Which of the following objectives would you like to prioritize: Efficiency, Soft switching, or Current stress?
3 To define operation conditions Current stress. Understood.Let’s discuss the operating conditions of your dual active bridge converter. Could you provide the following information: rated power, rated input voltage, rated output voltage, actual operating power, actual output voltage.
4 To select optimization algorithms Rated power: 1000 W Rated input voltage: 200 V Rated output voltage: 200 V Operating power: 600W Actual output voltage: 160V Thank you for providing the operating conditions. Lastly, we need to choose an optimization algorithm. Which optimization algorithm would you choose: PSO, DE or GA?
@@
PSO. OK, I have collected all the parameters and am about to execute the program.
@@
Table 5: An use case exemplar: with 5 request series, PHIA collects all necessary design specifications, which include a modulation strategy $S$ to design, prioritized objective $O$, operating conditions, and the selected optimization algorithms. After specification collection via text input, the design outcome is presented with text and figures.